# Gompertz Linear Units: Leveraging Asymmetry for Enhanced Learning Dynamics

### v1

@@ Appendix A Properties of GoLU: Further Details
Assuming the input distribution resembles a zero-centered, nearly-Gaussian form, which is likely particularly when employing batch normalization and appropriate weight initialization, the activations can be approximated by their tangents at the origin. Therefore a reduced slope at the origin translates into decreased sensitivity to input variations and lower output variance. We note that GoLU exhibits a lower slope magnitude not only in a neighborhood around the origin but across a significant portion of the negative input domain as illustrated in Figure 17.
Figure 17: Derivatives of the gate functions.
@@
Appendix B Details of the loss landscape experiment
@@
where $W_{\mathrm{trained}}$ are the trained weights of the model and $\alpha$ and $\beta$ are scalar values that determine the perturbation magnitude and are chosen as $\alpha,\beta\in[-1,1]$. For each pair of values $(\alpha,\beta)$, we compute the loss using the perturbed weights $W_{\mathrm{perturbed}}$ on a batch of test data. We then repeat this for a grid of $(\alpha,\beta)$ values to create a 3D surface plot as shown in Figure 5.
Appendix C Learning Rate ablation
For various tasks, we conduct a focused search over the learning rate to determine whether the default setting represents the optimal value and to assess its impact on the performance of models trained with different activation functions. Figures 19 and 19 present heatmaps of test results for Semantic Segmentation and Diffusion tasks, comparing models trained with various activation functions across different learning rates. For these tasks, the default learning rate, highlighted by a black box, differs from the optimal learning rate, indicated by a green box. Notably, while GoLU achieves a test mIoU that is a close second to ReLU under the default learning rate, it outperforms all other activation functions when evaluated at the optimal learning rate, which is consistent across all activations.
@@ Appendix C Learning Rate ablation
Motivated by these results, we further investigate the impact of learning rate on image classification tasks where GoLU demonstrated superior performance compared to baseline activations. Figures 21 and 21 present heatmaps of test accuracies for ResNet-50 and ViT-B/32 on ImageNet-1k.
@@
Appendix D Critical Difference Analysis
@@
Appendix E Training and inference times
@@
Appendix F Experimental Details
This section outlines detailed information about the datasets and training pipelines used for the various tasks studied in this work.
F.1 Image Classification - ImageNet
@@
F.2 Image Classification - CIFAR-10
@@ F.5 Object Detection
F.6 Instance Segmentation
The MS-COCO dataset for instance segmentation uses the same train and test sets as those used for Object Detection. Additionally, it trains with the exact same configurations used for Faster R-CNN-FPN in the previous subsection F.5.
F.7 Denoising Diffusion Probabilistic Models
@@
Appendix G Case Study: Bayesian Learning Curve Extrapolation using Prior-data fitted Networks
@@ Results
Figure  23 presents the log-scores for the final models, utilizing all 7 activation functions at 5 different learning rates, averaged over 3 training runs. At the original and optimal learning rate of 0.0001, GoLU ranks 6th among the 7 activations. However, a closer examination reveals that the choice of activation function seems to have minimal impact, as the differences between GoLU and the best (ELU) and worst (Swish) activation are within a single standard error. The learning rate ablation shows that GoLU ranks first at the highest stable learning rate (0.001), supporting previous findings that GoLU thrives in the high learning rate regime.

### v2

@@ Appendix A Properties of GoLU: Further Details
Assuming the input distribution resembles a zero-centered, nearly-Gaussian form, which is likely particularly when employing batch normalization and appropriate weight initialization, the activations can be approximated by their tangents at the origin. Therefore a reduced slope at the origin translates into decreased sensitivity to input variations and lower output variance. We note that GoLU exhibits a lower slope magnitude not only in a neighborhood around the origin but across a significant portion of the negative input domain as illustrated in Figure 2 (Right).
More generally, it can be shown analytically that for a given activation function $f$ that is smooth in a neighborhood of its input mean, and for a sufficiently localized input distribution, the variance of the activation output is approximately proportional to the square of its slope evaluated at the mean input $\mu$, with the input variance $\sigma^{2}$ serving as the proportionality constant
$\text{Var}[f(x)]\approx f^{\prime}(\mu)^{2}\sigma^{2}$ (7)
This formally demonstrates how smaller activation slopes result in reduced output variance.
To derive this connection, we apply the definition of variance to a scalar activation function $f(x)$:
$\text{Var}[f(x)]=\mathbb{E}[(f(x)-\mathbb{E}[f(x)])^{2}]=\mathbb{E}[f(x)^{2}]-\mathbb{E}[f(x)]^{2}$ (8)
Expanding $f(x)$ and $f(x)^{2}$ in a Taylor series around the mean input $\mu$, gives:
$f(x)$ $=$ $f(\mu)+f^{\prime}(\mu)(x-\mu)+\frac{1}{2}f^{\prime\prime}(\mu)(x-\mu)^{2}+\cdots$ (9)
$f(x)^{2}$ $=$ $f(\mu)^{2}+2f(\mu)f^{\prime}(\mu)(x-\mu)+f^{\prime}(\mu)^{2}(x-\mu)^{2}+f(\mu)f^{\prime\prime}(\mu)(x-\mu)^{2}+\cdots$ (10)
Taking expectations and using $\mathbb{E}[(x-\mu)]=0$ and $\mathbb{E}[(x-\mu)^{2}]=\sigma^{2}$ leads to:
$\mathbb{E}[f(x)]$ $=$ $f(\mu)+\frac{1}{2}f^{\prime\prime}(\mu)\sigma^{2}+\cdots$ (11)
$\mathbb{E}[f(x)^{2}]$ $=$ $f(\mu)^{2}+f^{\prime}(\mu)^{2}\sigma^{2}+f(\mu)f^{\prime\prime}(\mu)\sigma^{2}+\cdots$ (12)
Substituting these into the definition of the variance, and simplifying while retaining only the leading-order term, we obtain:
$\text{Var}[f(x)]$ $=$ $(f(\mu)^{2}+f^{\prime}(\mu)^{2}\sigma^{2}+f(\mu)f^{\prime\prime}(\mu)\sigma^{2}+\cdots)-(f(\mu)+\frac{1}{2}f^{\prime\prime}(\mu)\sigma^{2}+\cdots)^{2}$ (13)
$=$ $(f(\mu)^{2}+f^{\prime}(\mu)^{2}\sigma^{2}+f(\mu)f^{\prime\prime}(\mu)\sigma^{2}+\cdots)-(f(\mu)^{2}+f(\mu)f^{\prime\prime}(\mu)\sigma^{2}+\cdots)\approx f^{\prime}(\mu)^{2}\sigma^{2}$ (14)
which completes the proof.
@@
Appendix B Flipped Mish: a new self-gated activation with right-leaning distribution
Throughout this work, we have emphasized the influence of right-skewed asymmetry in the underlying distribution associated with an activation on model performance. To further explore this property, we leverage the left-skewed distribution underlying Mish to construct a new self-gated activation function exhibiting right-skewed asymmetry. This is achieved by reflecting the Mish-associated distribution about the vertical axis. Specifically, denoting the Mish distribution by $D(x)$ and its corresponding gate function by $g(x)$, the gate function for the flipped distribution $\tilde{D}(x)=D(-x)$ is given by $\tilde{g}(x)=1-g(-x)$, as shown by the following derivation:
[…]
where in the third equation we have redefined the dummy integration variable $y\rightarrow-y$ and in the last equation we have used the fact that the integral of the distribution $D(y)$ over its entire domain is equal to $1$. The resulting activation function, which we refer to as Flipped Mish (FMish), is thus defined as:
$\mathrm{FMish}(x)=x(1-\tanh(\mathrm{softplus}(-x)))$ (17)
Figure 8 compares FMish with GoLU, including their respective gate functions and the associated distributions.
Figure 8: Comparison of activations (left), gate functions (middle), and associated distributions (right) for FMish and GoLU.
We evaluated FMish on ResNet18, ResNet50 and ViT-B/32 trained on ImageNet-1k. Remarkably, it outperformed all baseline activations except GoLU, achieving $70.73\pm 0.05$ for ResNet18, $76.20\pm 0.01$ for ResNet50 and $75.67\pm 0.04$ for ViT-B/32 (compare with results in Table 2). This outcome aligns with expectations, as the slope of FMish at the origin is 0.4, lower than that of sigmoid and Gaussian CDF (0.5) but slightly higher than GoLU (0.37). These results further highlight the significance of right-leaning asymmetry and the resulting variance reduction.
Furthermore, we note that the Flipped Mish distribution does not decay as rapidly as the Gumbel distribution for large negative inputs, which may also contribute to its performance.
Appendix C Details of the loss landscape experiment
@@
where $W_{\mathrm{trained}}$ are the trained weights of the model and $\alpha$ and $\beta$ are scalar values that determine the perturbation magnitude and are chosen as $\alpha,\beta\in[-1,1]$. For each pair of values $(\alpha,\beta)$, we compute the loss using the perturbed weights $W_{\mathrm{perturbed}}$ on the full test set of the CIFAR-10 dataset. We then repeat this for a grid of $(\alpha,\beta)$ values to create a 3D surface plot as shown in Figure 5.
To provide a more quantitative understanding of the loss landscapes in Figure 5, we have plotted the density functions of the loss values for each activation function and computed their variance. The results, shown in Figure 9, indicate that GoLU achieves both a lower average loss and smaller variance compared to other activations, consistent with the observations from the 3D plots in Figure 5.
Figure 9: Comparison of loss value distributions across the loss landscape. GoLU achieves both a lower loss mean and variance.
Appendix D Learning Rate ablation
For various tasks, we conduct a focused search over the learning rate to determine whether the default setting represents the optimal value and to assess its impact on the performance of models trained with different activation functions. Figures 11 and 11 present heatmaps of test results for Semantic Segmentation and Diffusion tasks, comparing models trained with various activation functions across different learning rates. Figures 13 and 13 show similar heatmaps for the Instance Segmentation task, reporting Box mAP and Mask mAP, respectively. For these tasks, the default learning rate, highlighted by a black box, differs from the optimal learning rate, indicated by a green box. Notably, while GoLU performs slightly below the best-performing activation under the default learning rate, it outperforms all other activation functions when evaluated at the optimal learning rate, which is consistent across all activations.
@@ Appendix D Learning Rate ablation
Figure 12: Test Box mAP - Mask R-CNN-PFN ResNet-50 on MS-COCO. The default learning rate is 0.02 which is colored in black and the best learning rate is 0.03 which is colored in green.
Figure 13: Test Mask mAP - Mask R-CNN-PFN ResNet-50 on MS-COCO. The default learning rate is 0.02 which is colored in black and the best learning rate is 0.03 which is colored in green.
Motivated by these results, we further investigate the impact of learning rate on image classification tasks where GoLU demonstrated superior performance compared to baseline activations. Figures 15 and 15 present heatmaps of test accuracies for ResNet-50 and ViT-B/32 on ImageNet-1k.
@@
Appendix E Critical Difference Analysis
@@
Appendix F Training and inference times
Table 9 reports relative training and inference times with respect to baseline activations for our trained models. On average, GoLU achieves training and inference speeds on par with default activation functions, while offering improved performance. This makes GoLU a practical and effective alternative for training deep learning models.
@@ Appendix F Training and inference times
Appendix G Experimental Details
This section outlines detailed information about the datasets and training pipelines used for the various tasks studied in this work. All experiments in this section were conducted on NVIDIA A100 GPUs, with an approximate total compute time of 112K GPU hours, except for TinyViT, which was executed on an NVIDIA H100 GPU with a total runtime of 455 GPU hours.
G.1 Image Classification - ImageNet
@@
G.2 Image Classification - CIFAR-10
@@ G.5 Object Detection
G.6 Instance Segmentation
The MS-COCO dataset for instance segmentation uses the same train and test sets as those used for Object Detection. Additionally, it trains with the exact same configurations used for Faster R-CNN-FPN in the previous subsection G.5.
G.7 Denoising Diffusion Probabilistic Models
@@
Appendix H Test Loss Curves
To provide a more comprehensive view of GoLU’s test performance over the course of training, this section presents test curves, including loss and task-specific metrics, comparing GoLU with ReLU and GELU, and illustrating how their performance evolves throughout training.
Appendix I Machine Translation
To further assess GoLU across diverse tasks, we evaluated its performance on machine translation using the WMT14 English–German benchmark, with approximately 4.5 million training pairs. Specifically, we trained Transformer-Big models using the Fairseq framework (Ott et al., 2019), comparing GoLU against baseline activations including ReLU, which is the default in this architecture. The architecture follows the standard configuration with 6 encoder and 6 decoder layers, 1024-dimensional embeddings, 16 attention heads, and a feed-forward hidden size of 4096. Models were trained with three different random seeds for 50 epochs using the Adam optimizer ($\beta_{1}=0.9$, $\beta_{1}=0.98$), an inverse square root learning rate schedule (base LR $=5\times 10^{-4}$, 4000 warm-up steps), label smoothing of 0.1, and gradient accumulation of 16 to simulate large-batch training. Evaluation was conducted using beam search with BLEU4 as the performance metric. All runs were executed on a single NVIDIA L40S GPU with a total runtime of roughly 1750 GPU hours. As shown in Table 10, GoLU outperforms standard activation functions in terms of mean BLEU4 score, which highlights its effectiveness in sequence modeling as well.
Table 10: Mean and standard error of BLEU4 scores for Transformer-Big on the WMT14 English–German translation task.
Activation Test BLEU4
ReLU 28.33±0.14
LeakyReLU 28.26±0.04
ELU 27.49±0.08
GELU 28.20±0.04
Swish 28.34±0.08
Mish 28.31±0.10
GoLU 28.44±0.15
Appendix J Case Study: Bayesian Learning Curve Extrapolation using Prior-data fitted Networks
@@ Results
Figure  26 presents the log-scores for the final models, utilizing all 7 activation functions at 5 different learning rates, averaged over 3 training runs. At the original and optimal learning rate of 0.0001, GoLU ranks 6th among the 7 activations. However, a closer examination reveals that the choice of activation function seems to have minimal impact, as the differences between GoLU and the best (ELU) and worst (Swish) activation are within a single standard error. The learning rate ablation shows that GoLU ranks first at the highest stable learning rate (0.001), supporting previous findings that GoLU thrives in the high learning rate regime.